ISS Daily Summary Report – 09/23/14

SpaceX-4 Dragon Capture, Berth, and Ingress: After a nominal Dragon rendezvous with ISS, the Crew successfully captured Dragon at 5:52am CDT utilizing the Space Station Remote Manipulator System (SSRMS).  The Dragon capsule was subsequently berthed to the Node 2 Nadir Port at 8:15am CDT. Once berthed, the Crew completed pressurization and outfitting of the Node 2 to Dragon vestibule.  With a few hours remaining at the end of the day, the Crew proceeded with ingress activities which were originally planned for tomorrow. Dragon hatch was opened today at 12:52pm CDT. The Crew will pick up with Dragon Cargo transfer operations tomorrow.

Circadian Rhythms:  Wiseman set-up instrumentation and configured sensors to support the Flight Day 120 36 hour continuous Circadian rhythm session.  After the measurement, de-instrument, data transfer, and stow was performed.  Circadian Rhythms investigates the role of synchronized circadian rhythms, or the “biological clock,” and how it changes during long-duration spaceflight. Researchers hypothesize that a non-24-hour cycle of light and dark affects crewmembers’ circadian clocks. The investigation also addresses the effects of reduced physical activity, microgravity and an artificially controlled environment. Changes in body composition and body temperature, which also occur in microgravity, can affect crewmembers’ circadian rhythms as well. Understanding how these phenomena affect the biological clock will improve performance and health for future crewmembers.

Rodent Research Transfer Review:  Wiseman and Gerst reviewed Rodent Research Animal Transfer material in preparation for transferring the animals from the Transporter to the Habitat.  This activity was delayed one day due to delayed berthing of SpaceX-4 vehicle. 

Treadmill 2 (T2) Issue: Yesterday, the crew was having issues remaining logged into the Station Support Computer (SSC) connected to T2.  After eventually connecting, the crew found the treadmill’s belt would not move, although indications from the computer appeared to show it was operating as expected.  Attempts to regain functionality from the ground using the remote troubleshooting procedure were unsuccessful.  T2 Rack and SSC power cycles were also unsuccessful.   An attempt to utilize the redundant Ethernet port on the T2 rack did not resolve the issue. This morning the crew was asked to install a debug cable to the SSC.  Having this cable installed will allow ground teams to continue troubleshooting with increased insight into the system.  Currently, T2 is not operational and teams are defining the next series of troubleshooting steps.

Ground Controlled Mobile Servicing System (MSS) Operations:  Overnight, the Robotics Ground Controllers powered up the MSS in the Hot Backup Configuration and used a SSRMS camera to survey the Node 2 Nadir Active Common Berthing Mechanism (CBM) sealing surface.  After the crew captured Dragon, the Robotics Ground Controllers used the SSRMS to maneuver Dragon into position for the inspection of the Passive CBM and then maneuvered Dragon for installation on the Node-2 Nadir Active CBM at approximately 8:15am CDT.

Today’s Planned Activities

All activities have been completed unless otherwise noted.

  • CCP panel – cable routing and testing
  • Laptop RS1(2) & RSS2 reboot / r/g 1009 step 1
  • SM Caution and Warning Panel test (33)
  • USOS window shutters close
  • WRS – water sample analysis
  • RWS config to observe DRAGON rendezvous
  • URAGAN. Observation and photography / r/g 6683
  • Camera install to feed video to LAB RWS monitor 3
  • Cupola Minicam config to support Dragon docking
  • Intravehicular Activity (IVA) Clothing
  • DRAGON. Arriving to R-bar
  • S/G1 comm check with Baikonur
  • KONTROL. Switching of “Индикатор-МКС” science hardware measurement modes./ r/g 6684
  • Dragon – arriving to R-bar
  • ПФ1, ПФ2 dust filters replacement and В1,B2 fan grill cleaning in MRM2
  • Location preparation for ТБУ-В thermostat install / r/g 6687
  • Dragon – SSRMS capture
  • CBCS – camera install on Node 2 nadir hatch and power up
  • TOCA – data recording
  • IVA – exercise clothing questionnaire
  • CBM – capture and bolt driving
  • Flushing В1 connector on БВ1 Rodnik tank of ТКГ423 (DC1) / r/g 6681
  • CBCS – removing camera from Node 2 nadir hatch
  • ABOUT GAGARIN FROM SPACE – Ham radio session with Equinoccial Technological University in Kito
  • Deconfigure cupola from rendzvs operations
  • СОЖ maintenance
  • CCP – cable routing
  • Dragon – repress from Node 2 and leak check
  • ИП-1 sensors monitoring / СКДС п.3.9 стр. 3-5 (14)
  • IVA – stowing dried clothes
  • VISIR – experiment run. Tagup  / r/g 6685
  • Dragon – vestibule outfitting
  • Electron-VM БКО R&R/ r/g 6682
  • Node 2 – Removing docking bolts control unit
  • RR- hw transfer
  • Thermolab – circadian rhythms equipment ops
  • IVA – everyday clothing questionnaire

Completed Task List Items

  • ARED EVAC [Completed Sunday GMT 264]
  • RPOP Setup  [Complete]
  • Replace KTO [Complete]
  • FPEF cable connect [In Work]
  • T2 Troubleshooting [Complete]
  • LAB1OP4 General Luminaire Assembly (GLA) R&R  [Complete]


  • System commanding associated with Dragon Rendezvous and Berth [Complete]

Three-Day Look Ahead:

  • Wednesday, 09/24: T2 SSC Swap, Dragon Unpack, Rodent Research Transfer, Micro8 Transfer
  • Thursday, 09/25: 40S Launch and Dock, Safety Briefing, Zebra Fish Water Exchange
  • Friday, 09/26: Crew Off Duty

QUICK ISS Status – Environmental Control Group:







[СКВ] 1 – SM Air Conditioner System (“SKV1”)


[СКВ] 2 – SM Air Conditioner System (“SKV2”)


Carbon Dioxide Removal Assembly (CDRA) Lab


Carbon Dioxide Removal Assembly (CDRA) Node 3


Major Constituent Analyzer (MCA) Lab


Major Constituent Analyzer (MCA) Node 3


Oxygen Generation Assembly (OGA)


Urine Processing Assembly (UPA)


Trace Contaminant Control System (TCCS) Lab


Trace Contaminant Control System (TCCS) Node 3

Full up